# State of the Art in Text Classification for South Slavic Languages: Fine-Tuning or Prompting?

### v1

@@ Abstract
Until recently, fine-tuned BERT-like models provided state-of-the-art performance on text classification tasks. With the rise of instruction-tuned decoder-only models, commonly known as large language models (LLMs), the field has increasingly moved toward zero-shot and few-shot prompting. However, the performance of LLMs on text classification, particularly on less-resourced languages, remains under-explored. In this paper, we evaluate the performance of current language models on text classification tasks across several South Slavic languages. We compare openly available fine-tuned BERT-like models with a selection of open-source and closed-source LLMs across three tasks in three domains: sentiment classification in parliamentary speeches, topic classification in news articles and parliamentary speeches, and genre identification in web texts. Our results show that LLMs demonstrate strong zero-shot performance, often matching or surpassing fine-tuned BERT-like models. Moreover, when used in a zero-shot setup, LLMs perform comparably in South Slavic languages and English. However, we also point out key drawbacks of LLMs, including less predictable outputs, significantly slower inference, and higher computational costs. Due to these limitations, fine-tuned BERT-like models remain a more practical choice for large-scale automatic text annotation.
@@ 1.  Introduction
Until recently, the dominant approach for text classification tasks relied on fine-tuning BERT-like transformer models on thousands of manually-annotated training examples. Recently, however, the field has shifted with the development of instruction-tuned decoder-only transformer models. These models, also commonly referred to as large language models (LLMs), which were originally developed primarily for text generation tasks, have demonstrated remarkable capabilities across a broad range of natural language processing (NLP) tasks, including text classification (kuzman2023automatic; huang2023chatgpt).
In this paper, we focus on South Slavic languages, where research on text classification tasks included in our study has, until recently, been limited or even non-existent (kuzman2023survey; mochtak2024parlasent; kuzman-iptc-classification). We take a first step toward systematically evaluating the current state of the art for text classification in these languages. Our evaluation is based on three text classification tasks in three different domains for which manually-annotated test datasets in South Slavic languages and fine-tuned BERT-like classifiers are freely available: sentiment classification of parliamentary speeches, topic classification in news articles, topic classification in parliamentary speeches, and automatic genre identification in web texts. These tasks span different domains and language styles, allowing for a comprehensive analysis of the performance of transformer-based models on text classification tasks. Specifically, we compare the performance of openly available fine-tuned BERT-like models with the zero-shot capabilities of both open-source and closed-source LLMs used via prompting.
@@
By evaluating various models on a selection of text classification tasks in English and various South Slavic languages, we set out to test the following two hypotheses that are based on previous experiments with fine-tuned BERT-like models and LLMs on automatic genre identification (kuzman2023automatic), news topic classification (kuzman-iptc-classification) and sentiment analysis in parliamentary texts (mochtak2025parlasent):
@@ 2.  Related Work
After the introduction of transformer architectures, BERT (bidirectional encoder representations from transformers) models have achieved state-of-the-art results in text classification tasks, outperforming earlier non-neural approaches, such as support vector machines (SVMs). They have also demonstrated strong cross-lingual zero-shot capabilities in various classification tasks, including automatic genre identification (kuzman2023survey), news topic classification (petukhova2023mn; de2020news), and sentiment classification (mochtak2024parlasent). However, these models still require fine-tuning on a training dataset, developed during manual annotation campaigns that are time-consuming and costly.
Instruction-tuned decoder-only transformer models, commonly referred to as large language models (LLMs), have recently shown strong performance in a range of classification tasks, even in zero-shot prompting setups that require no training data (kuzman2023automatic; ljubevsic2024dialect; huang2023chatgpt). They have achieved promising results on various natural language processing tasks, including stance detection (zhang2022would), implicit hate speech categorization (huang2023chatgpt), news topic classification (kuzman-iptc-classification), automatic genre identification (kuzman2023automatic), causal commonsense reasoning (ljubevsic2024jsi), and machine translation (hendy2023good). Due to their promising performance, researchers have even started using them as data annotators, either by generating text and labels (meng2022generating) or by annotating pre-existing texts (kuzman-iptc-classification). Despite the growing interest in this topic, the majority of evaluations of LLMs used in text classification tasks are limited only to English (sun2023text; zhang2025pushing; kostina2025large; zhao2024advancing). Systematic multilingual evaluations, especially which would include less-resourced languages such as those in the South Slavic group remain limited. Our work addresses this gap by providing a comparative evaluation of open-source and closed-source LLMs with openly-available fine-tuned BERT-like models across four benchmark families comprising three diverse classification tasks and three different domains in South Slavic languages and English.
@@ 3.  Benchmarks
The topic classification task is evaluated on two domains: 1) news articles, namely, the Croatian and Slovenian IPTC test datasets (kuzman-iptc-classification), which comprise around 300 text instances per language, and 2) parliamentary speeches, namely, the Bosnian, Croatian, English and Serbian ParlaCAP test datasets that consist of approximately 820 to 880 instances per language. In the ParlaCAP benchmarks, an instance is a transcription of an utterance given by a parliamentary member in a parliamentary session.
The topic classification task involves the highest number of labels, that is, 17 news topic labels from the top level of the IPTC NewsCodes Media Topic hierarchical schema [^1] https://show.newscodes.org/index.html?newscodes=medtop&lang=en-GB&startTo=Show (iptcGroupsNewsCodes), and 22 agenda setting topic labels (21 major topics and a label Other) from the Comparative Agendas Project (CAP) (baumgartner2019comparative) Master Codebook (bevan2019gone). [^2] https://www.comparativeagendas.net/pages/master-codebook
In contrast, the Bosnian, Croatian, English, and Serbian ParlaSent sentiment identification datasets (mochtak2024parlasent; parlasent-repository) have a significantly lower granularity of labels, with only 3 categories. They are represented by the largest number of instances, ranging from 190 (Bosnian part) to 2600 (English part) sentence-level instances.
With 8 labels, the Croatian, English, Macedonian, and Slovenian GINCO genre datasets (kuzman2023automatic) represent a midpoint in label granularity among the four benchmark families. However, the genre identification task might be the most difficult one, as genre identification depends on the interpretation of full texts with the focus on author’s purpose, the common function of the text, and the text’s conventional form (orlikowski1994genre). This complexity has also contributed to smaller test datasets in terms of the number of text instances, as manual annotation is more time-consuming. It is also important to note that, unlike the parliamentary datasets, the English portion of the genre datasets is not fully comparable to the South Slavic portions, which are label-balanced and contain fewer ambiguous instances. Nevertheless, the genre datasets remain valuable for evaluating model performance within each language.
All test datasets were manually annotated by annotators that are deemed reliable based on their satisfactory inter-annotator agreement, namely, Krippendorff’s alpha (krippendorff2018content) values close to or above the 0.667 threshold for reliable annotation. To prevent large language models from incorporating the test datasets during their training phase, the test datasets are not publicly available, except for the ParlaSent benchmark family. Access to other datasets is granted on request from the corresponding authors. Further details on the test datasets are provided in Section A.1 of the Appendix.
@@ 4.  Methodology
In this paper, we evaluate the main machine learning approaches that have recently been used for our selection of text classification tasks, with the focus on the comparison between the freely available fine-tuned BERT models and the open-source and closed-source LLMs. [^3] The code for the model evaluation and analysis of results is available at https://github.com/TajaKuzman/Benchmarking-Text-Classification-on-South-Slavic. The models are evaluated on four families of test datasets that comprise South Slavic languages. The performance of the models is evaluated based on the micro-F1 and macro-F1 metrics, which enable assessment of the model performance at both the instance and label levels, respectively.
@@
  • fine-tuned BERT-like classifiers: in our study, we evaluate previously developed openly accessible multilingual fine-tuned BERT-like models that have been fine-tuned for the respective task, namely, the XLM-R-ParlaSent (parlasent-model; mochtak2024parlasent) model for sentiment identification in parliamentary texts, the X-GENRE classifier (kuzman2023automatic; x-genre-repository; x-genre-classifier-huggingface) for automatic genre identification, the IPTC News Topic classifier (kuzman-iptc-classification; iptc_model_huggingface) for news topic classification, and the ParlaCAP classifier (parlacap_model) for topic classification in parliamentary speeches. The XLM-R-ParlaSent and the ParlaCAP models are based on the XLM-R-parla pretrained model (xlm-r-parla) that was developed by additionally pretraining the large-sized XLM-RoBERTa model (conneau2020unsupervised) on parliamentary proceedings in 30 European languages (mochtak2024parlasent). The XLM-R-ParlaSent model was fine-tuned on 13 thousand instances from the ParlaSent sentiment training dataset (parlasent-repository) in seven European languages (Bosnian, Croatian, Czech, English, Serbian, Slovak, and Slovenian) (mochtak2024parlasent), while the ParlaCAP model was fine-tuned on around 30 thousand speeches from parliamentary debates annotated with CAP topic labels, originating from the ParlaMint 4.1 parliamentary datasets (parlamint_41; erjavec2025parlamint) from 29 European parliaments. The X-GENRE classifier is based on the base-sized XLM-RoBERTa model (conneau2020unsupervised) and was fine-tuned on the training split of the X-GENRE dataset (x-genre-dataset) in English and Slovenian; while the IPTC News Topic classifier is based on the large-sized XLM-RoBERTa model (conneau2020unsupervised) that was fine-tuned on the EMMediaTopic dataset (emmediatopic) in Catalan, Croatian, Greek, and Slovenian. All fine-tuned models use the same classes as the test datasets used in our study.
@@
  • open-source and closed-source large language models: we use closed-source OpenAI models, namely the GPT-3.5-Turbo (gpt-3.5-turbo-0125) (openai), GPT-4o (gpt-4o-2024-08-06) (openai-gpt4o) and the GPT-5 (gpt-5-2025-08-07) (gpt-5); a closed-source Gemini 2.5 Flash model (comanici2025gemini) by Google DeepMind; a closed-source Mistral Medium 3.1 model (mistral-medium-2508) (mistral) by Mistral AI; and four open-source models, namely, the Meta LLaMA 3.3 model (llama33modelcard), the Gemma 3 model (gemma3), the Qwen 3 model (yang2025qwen3), and the DeepSeek-R1-Distill model (DeepSeek-R1-Distill-Qwen-14B) (guo2025deepseek). It is important to note that while the LLaMA model was pretrained on a web text collection in various languages, it is said to support only 8 languages, namely English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai (llama33modelcard). The DeepSeek-R1-Distill model is based on the Qwen 2.5 model (qwen2.5; team2024qwen2) that provides support for more than 29 languages – not including South Slavic languages though. In contrast, the Gemma 3 model is reported to support over 140 languages (gemma3), and the Qwen 3 model was pretrained on 119 languages (yang2025qwen3). While closed-source models are said to be massively multilingual, with Gemini 2.5 models being pretrained on over 400 languages (comanici2025gemini), details on their language coverage are very limited.
@@
Open-source models were installed locally and executed via the Ollama API service (maric_local_llms). OpenAI models were used through the chat completion endpoint via the OpenAI API, whereas other closed-source models were accessed through the OpenRouter platform [^4] https://openrouter.ai/ that provides a unified API access to various closed-source models. To prevent any bias, all models were used with their default parameters. The only parameter that we defined is the temperature which we set to 0 to ensure a more deterministic behaviour of the models. More details on the models and their implementation, including information on the availability of openly available models and fine-tuning datasets, are provided in Section A.2 of the Appendix.
@@
All instruction-tuned LLMs are used in a zero-shot prompting setup, meaning that they receive only a task description and label definitions. The models are instructed to output a label, represented by a digit. The same prompt per benchmark family is used for all LLMs. Prompts are provided in Figure 4 in Section A.2 of the Appendix.
@@ 5.  Results
In this section, we evaluate the performance of fine-tuned BERT-like models and the instruction-tuned LLMs on a selection of text classification tasks that include test datasets in South Slavic languages. First, in Section 5.1, we provide results on the four benchmark families with the focus on hypothesis H1, which expects that zero-shot prompting with LLMs can provide performance that is comparable to that of fine-tuned BERT-like models. In Section 5.2, we compare in more detail the performance of the closed-source and open-source LLMs on the three text classification tasks, which is followed by a discussion on the advantages and limitations of LLMs for data annotation based on text classification tasks (Section 5.3). Lastly, in Section 5.4, we compare the performance of LLMs on English test datasets with their performance on South Slavic datasets, addressing hypothesis H2, which presumes that the available multilingual LLMs perform similarly on South Slavic languages as on English.

### v2

@@ Abstract
Until recently, fine-tuned BERT-like models provided state-of-the-art performance on text classification tasks. With the rise of instruction-tuned decoder-only models, commonly known as large language models (LLMs), the field has increasingly moved toward zero-shot and few-shot prompting. However, the performance of LLMs on text classification, particularly on less-resourced languages, remains under-explored. In this paper, we evaluate the performance of current language models on text classification tasks across several South Slavic languages. We compare openly available fine-tuned BERT-like models with a selection of open-weight and closed-source LLMs across three tasks in three domains: sentiment classification in parliamentary speeches, topic classification in news articles and parliamentary speeches, and genre identification in web texts. Our results show that LLMs demonstrate strong zero-shot performance, often matching or surpassing fine-tuned BERT-like models. Moreover, when used in a zero-shot setup, LLMs perform comparably in South Slavic languages and English. However, we also point out key drawbacks of LLMs, including less predictable outputs, significantly slower inference, and higher computational costs. Due to these limitations, fine-tuned BERT-like models remain a more practical choice for large-scale automatic text annotation.
@@ 1.  Introduction
Until recently, the dominant approach for text classification tasks relied on fine-tuning BERT-like transformer models on thousands of manually-annotated training examples. Recently, however, the field has shifted with the development of instruction-tuned decoder-only transformer models. These models, also commonly referred to as large language models (LLMs), which were originally developed primarily for text generation tasks, have demonstrated remarkable capabilities across a broad range of natural language processing (NLP) tasks, including text classification (Kuzman et al., 2023; Huang et al., 2023).
In this paper, we focus on South Slavic languages, where research on text classification tasks included in our study has, until recently, been limited or even non-existent (Kuzman and Ljubešić, 2023; Mochtak et al., 2024; Kuzman and Ljubešić, 2025). We take a first step toward systematically evaluating the current state of the art for text classification in these languages. Our evaluation is based on three text classification tasks in three different domains for which manually-annotated test datasets in South Slavic languages and fine-tuned BERT-like classifiers are freely available: sentiment classification of parliamentary speeches, topic classification in news articles, topic classification in parliamentary speeches, and automatic genre identification in web texts. These tasks span different domains and language styles, allowing for a comprehensive analysis of the performance of transformer-based models on text classification tasks. Specifically, we compare the performance of openly available fine-tuned BERT-like models with the zero-shot capabilities of both open-weight and closed-source LLMs used via prompting.
@@
By evaluating various models on a selection of text classification tasks in English and various South Slavic languages, we set out to test the following two hypotheses that are based on previous experiments with fine-tuned BERT-like models and LLMs on automatic genre identification (Kuzman et al., 2023), news topic classification (Kuzman and Ljubešić, 2025) and sentiment analysis in parliamentary texts (Mochtak et al., 2025):
@@ 2.  Related Work
After the introduction of transformer architectures, BERT (bidirectional encoder representations from transformers) models have achieved state-of-the-art results in text classification tasks, outperforming earlier non-neural approaches, such as support vector machines (SVMs). They have also demonstrated strong cross-lingual zero-shot capabilities in various classification tasks, including automatic genre identification (Kuzman and Ljubešić, 2023), news topic classification (Petukhova and Fachada, 2023; De Clercq et al., 2020), and sentiment classification (Mochtak et al., 2024). However, these models still require fine-tuning on a training dataset, developed during manual annotation campaigns that are time-consuming and costly.
Instruction-tuned decoder-only transformer models, commonly referred to as large language models (LLMs), have recently shown strong performance in a range of classification tasks, even in zero-shot prompting setups that require no training data (Kuzman et al., 2023; Ljubešić et al., 2024a; Huang et al., 2023; Kuzman Pungeršek et al., 2026). They have achieved promising results on various natural language processing tasks, including stance detection (Zhang et al., 2022), implicit hate speech categorization (Huang et al., 2023), news topic classification (Kuzman and Ljubešić, 2025), automatic genre identification (Kuzman et al., 2023), causal commonsense reasoning (Ljubešić et al., 2024b), and machine translation (Hendy et al., 2023). Due to their promising performance, researchers have even started using them as data annotators, either by generating text and labels (Meng et al., 2022) or by annotating pre-existing texts (Kuzman and Ljubešić, 2025; Kuzman Pungeršek et al., 2026). Despite the growing interest in this topic, the majority of evaluations of LLMs used in text classification tasks are limited only to English (Sun et al., 2023; Zhang et al., 2025; Kostina et al., 2025; Zhao et al., 2024). Systematic multilingual evaluations, especially which would include less-resourced languages such as those in the South Slavic group remain limited. Our work addresses this gap by providing a comparative evaluation of open-weight and closed-source LLMs with openly-available fine-tuned BERT-like models across four benchmark families comprising three diverse classification tasks and three different domains in South Slavic languages and English.
@@ 3.  Benchmarks
The topic classification task is evaluated on two domains: 1) news articles, namely, the Croatian and Slovenian IPTC test datasets (Kuzman and Ljubešić, 2025), which comprise around 300 text instances per language, and 2) parliamentary speeches, namely, the Bosnian, Croatian, English and Serbian ParlaCAP test datasets (Kuzman Pungeršek et al., 2026) that consist of approximately 820 to 880 instances per language. In the ParlaCAP benchmarks, an instance is a transcription of an utterance given by a parliamentary member in a parliamentary session.
The topic classification task involves the highest number of labels, that is, 17 news topic labels from the top level of the IPTC NewsCodes Media Topic hierarchical schema [^1] https://show.newscodes.org/index.html?newscodes=medtop&lang=en-GB&startTo=Show (IPTC, 2022), and 22 policy topic labels (21 major topics and a label Other) from the Comparative Agendas Project (CAP; Baumgartner et al., 2019) Master Codebook (Bevan, 2019). [^2] https://www.comparativeagendas.net/pages/master-codebook
In contrast, the Bosnian, Croatian, English, and Serbian ParlaSent sentiment identification datasets (Mochtak et al., 2024; Mochtak et al., 2023) have a significantly lower granularity of labels, with only 3 categories. They are represented by the largest number of instances, ranging from 190 (Bosnian part) to 2600 (English part) sentence-level instances.
With 8 labels, the Croatian, English, Macedonian, and Slovenian GINCO genre datasets (Kuzman et al., 2023) represent a midpoint in label granularity among the four benchmark families. However, the genre identification task might be the most difficult one, as genre identification depends on the interpretation of full texts with the focus on author’s purpose, the common function of the text, and the text’s conventional form (Orlikowski and Yates, 1994). This complexity has also contributed to smaller test datasets in terms of the number of text instances, as manual annotation is more time-consuming. It is also important to note that, unlike the parliamentary datasets, the English portion of the genre datasets is not fully comparable to the South Slavic portions, which are label-balanced and contain fewer ambiguous instances. Nevertheless, the genre datasets remain valuable for evaluating model performance within each language.
All test datasets were manually annotated by annotators that are deemed reliable based on their satisfactory inter-annotator agreement, namely, Krippendorff’s alpha (Krippendorff, 2018) values close to or above the 0.667 threshold for reliable annotation. To prevent large language models from incorporating the test datasets during their training phase, the test datasets are not publicly available, except for the ParlaSent benchmark family. Access to other datasets is granted on request from the corresponding authors. Further details on the test datasets are provided in Section A.1 of the Appendix.
@@ 4.  Methodology
In this paper, we evaluate the main machine learning approaches that have recently been used for our selection of text classification tasks, with the focus on the comparison between the freely available fine-tuned BERT models and the open-weight and closed-source LLMs. [^3] The code for the model evaluation and analysis of results is available at https://github.com/TajaKuzman/Benchmarking-Text-Classification-on-South-Slavic. The models are evaluated on four families of test datasets that comprise South Slavic languages. The performance of the models is evaluated based on the micro-F1 and macro-F1 metrics, which enable assessment of the model performance at both the instance and label levels, respectively.
@@
  • fine-tuned BERT-like classifiers: in our study, we evaluate previously developed openly accessible multilingual fine-tuned BERT-like models that have been fine-tuned for the respective task, namely, the XLM-R-ParlaSent (Rupnik et al., 2023; Mochtak et al., 2024) model for sentiment identification in parliamentary texts, the X-GENRE classifier (Kuzman et al., 2023; Kuzman and Ljubešić, 2024d, 2023) for automatic genre identification, the IPTC News Topic classifier (Kuzman and Ljubešić, 2025; Kuzman and Ljubešić, 2025) for news topic classification, and the ParlaCAP classifier (Kuzman Pungeršek et al., 2026; Kuzman Pungeršek and Ljubešić, 2025) for topic classification in parliamentary speeches. The XLM-R-ParlaSent and the ParlaCAP models are based on the XLM-R-parla pretrained model (Ljubešić et al., 2023) that was developed by additionally pretraining the large-sized XLM-RoBERTa model (Conneau et al., 2020) on parliamentary proceedings in 30 European languages (Mochtak et al., 2024). The XLM-R-ParlaSent model was fine-tuned on 13 thousand instances from the ParlaSent sentiment training dataset (Mochtak et al., 2023) in seven European languages (Bosnian, Croatian, Czech, English, Serbian, Slovak, and Slovenian; Mochtak et al., 2024), while the ParlaCAP model was fine-tuned on the ParlaCAP-train dataset (Kuzman Pungeršek and Ljubešić, 2026; Kuzman Pungeršek et al., 2026) that comprises around 30 thousand speeches from parliamentary debates annotated with CAP topic labels, originating from the ParlaMint 4.1 parliamentary datasets (Erjavec et al., 2024; Erjavec et al., 2025) in 29 European languages. The X-GENRE classifier is based on the base-sized XLM-RoBERTa model (Conneau et al., 2020) and was fine-tuned on the training split of the X-GENRE dataset (Kuzman and Ljubešić, 2024a) in English and Slovenian; while the IPTC News Topic classifier is based on the large-sized XLM-RoBERTa model (Conneau et al., 2020) that was fine-tuned on the EMMediaTopic dataset (Kuzman and Ljubešić, 2024c) in Catalan, Croatian, Greek, and Slovenian. All fine-tuned models use the same classes as the test datasets used in our study.
@@
  • open-weight and closed-source large language models: we use closed-source OpenAI models, namely the GPT-3.5-Turbo (gpt-3.5-turbo-0125; OpenAI, 2023), GPT-4o (gpt-4o-2024-08-06; OpenAI, 2024) and the GPT-5 (gpt-5-2025-08-07; OpenAI, 2025); a closed-source Gemini 2.5 Flash model (Comanici et al., 2025) by Google DeepMind; a closed-source Mistral Medium 3.1 model (mistral-medium-2508; Mistral AI, 2025) by Mistral AI; and four open-weight models, namely, the Meta LLaMA 3.3 model (Meta, 2024), the Gemma 3 model (Gemma Team et al., 2025), the Qwen 3 model (Yang et al., 2025), and the DeepSeek-R1-Distill model (DeepSeek-R1-Distill-Qwen-14B; Guo et al., 2025). It is important to note that while the LLaMA model was pretrained on a web text collection in various languages, it is said to support only 8 languages, namely English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai (Meta, 2024). The DeepSeek-R1-Distill model is based on the Qwen 2.5 model (Qwen Team, 2024b, a) that provides support for more than 29 languages – not including South Slavic languages though. In contrast, the Gemma 3 model is reported to support over 140 languages (Gemma Team et al., 2025), and the Qwen 3 model was pretrained on 119 languages (Yang et al., 2025). While closed-source models are said to be massively multilingual, with Gemini 2.5 models being pretrained on over 400 languages (Comanici et al., 2025), details on their language coverage are very limited.
@@
Open-weight models were installed locally and executed via the Ollama API service (Marić et al., 2025). OpenAI models were used through the chat completion endpoint via the OpenAI API, whereas other closed-source models were accessed through the OpenRouter platform [^4] https://openrouter.ai/ that provides a unified API access to various closed-source models. To prevent any bias, all models were used with their default parameters. The only parameter that we defined is the temperature which we set to 0 to ensure a more deterministic behaviour of the models. More details on the models and their implementation, including information on the availability of openly available models and fine-tuning datasets, are provided in Section A.2 of the Appendix.
@@
All instruction-tuned LLMs are used in a zero-shot prompting setup, meaning that they receive only a task description and label definitions. All prompts and label definitions are written in English, while the instances are provided in the original languages (English or South Slavic). Changing the language of the prompt could introduce an additional factor that affects performance. In the presented experiments, our goal is to assess model performance on a specific task and language, rather than to evaluate their instruction-following abilities across different languages. The models are instructed to output a label, represented by a digit. The same prompt per benchmark family is used for all LLMs. Prompts are provided in Figure 4 in Section A.2 of the Appendix.
@@ 5.  Results
In this section, we evaluate the performance of the fine-tuned BERT-like models and the instruction-tuned LLMs on a selection of text classification tasks that include test datasets in South Slavic languages. First, in Section 5.1, we provide results on the four benchmark families with the focus on hypothesis H1, which expects that zero-shot prompting with LLMs can provide performance that is comparable to that of fine-tuned BERT-like models. In Section 5.2, we compare in more detail the performance of the closed-source and open-weight LLMs on the three text classification tasks, which is followed by a discussion on the advantages and limitations of LLMs for data annotation based on text classification tasks (Section 5.3). Lastly, in Section 5.4, we compare the performance of LLMs on English test datasets with their performance on South Slavic datasets, addressing hypothesis H2, which presumes that the available multilingual LLMs perform similarly on South Slavic languages as on English.